OpenStudy (anonymous):

What is the slope of the line that passes through the points (-2, 15) and (-8, -5)?

OpenStudy (anonymous):

The slope is given by (change in y)/(change in x) You are given two coordinates of the form (x,y) So the slope between them is given by; dy/dx=(15--5)/(-2--8)=(15+5)/(6)=20/6=10/3